# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

### v2

@@ Abstract
Foundation models trained on internet-scale data, such as Vision-Language Models (VLMs), excel at performing a wide variety of common sense tasks like visual question answering. Despite their impressive capabilities, these models cannot currently be directly applied to challenging robot manipulation problems that require complex and precise continuous reasoning over long horizons. Task and Motion Planning (TAMP) systems can control high-dimensional continuous systems over long horizons via a hybrid search over traditional primitive robot skills. However, these systems require detailed models of how the robot can impact its environment, preventing them from directly interpreting and addressing novel human objectives, for example, an arbitrary natural language goal. We propose deploying VLMs within TAMP systems by having them generate discrete and continuous language-parameterized constraints that enable TAMP to reason about open-world concepts. Specifically, we propose algorithms for VLM partial planning that constrain a TAMP system’s discrete temporal search and VLM continuous constraints interpretation to augment the traditional manipulation constraints that TAMP systems seek to satisfy. Experiments demonstrate that our approach — OWL-TAMP — outperforms several related baselines, including those that solely use TAMP or VLMs for planning, across several long-horizon manipulation tasks specified directly through natural language. We additionally demonstrate that our approach is compatible with a variety of TAMP systems and can be deployed to solve challenging manipulation tasks on real-world hardware. Project website: https://owl-tamp.csail.mit.edu/. Work partially conducted at NVIDIA Research. Correspondence to njk@csail.mit.edu, cgarrett@nvidia.com.
I Introduction
The advent of foundation models trained on internet-scale data has led to unprecedented progress on traditionally-hard tasks in vision and natural language. Current Large Language Models (LLMs) and Vision-Language Models (VLMs) are able to complete text from partial specifications, answer questions about images, and even solve challenging word problems that require reasoning and common sense [1, 2, 3]. This impressive performance has inspired several systems that attempt to use existing pretrained models in robotics [4, 5, 6, 7]. Such systems exhibit impressive flexibility: unlike classical robotics approaches, they are able to accomplish novel goals specified by natural language or images. However, currently no publicly-available foundation models exist that can directly output continuous values (e.g. joint angles, grasps, placements), which are critical for interacting with the physical world.
In contrast, classical Task and Motion Planning (TAMP) systems are capable of solving complex and long-horizon tasks ranging from setting a dining table to 3D printing complex structures [8, 9, 10]. These systems leverage models of the robot and its environment to explicitly reason about both discrete and continuous values in robotics problems. While such systems are powerful on the set of problems they have been designed for, they do not transfer to novel problems for which their models are unspecified. Enabling a TAMP system to solve novel problems often requires manually extending the underlying model, which is tedious and not scalable when operating in unstructured human environments.
We are interested in combining the complementary benefits of foundation-models and TAMP to tackle long-horizon manipulation tasks that are open world, namely where the vocabulary of objectives is unbounded. Specifically, we assume tasks are specified in natural language, which may involve concepts that an underlying TAMP system does not have built-in, but which can be achieved by chaining together robot motion primitives the TAMP system possesses. As an example, a TAMP system that is capable of accomplishing pick-and-place tasks expects goals in the form of logical expressions involving predicates like On(apple, plate). Consider the goal in Figure LABEL:fig:system: “Put the orange on the table where the apple initially is”. This goal cannot be expressed in terms of On, and thus there would be no way a TAMP system could solve it, even though it could be accomplished by a sequence of pick-place primitives. A pure VLM system would also struggle with this task since it must not only predict that the apple needs to be moved out of the way before the orange can be placed, but also continuous robot motions that realize this.
Our key insight is that we can integrate the discrete-continuous planning of TAMP systems and common sense reasoning of VLMs through the contract of constraints. In particular, VLMs are capable of mapping a very wide range of open world expressions into discrete action sequences (e.g. that a potato must be cooked before it can be served), and code that represents continuous constraints over important decision variables (e.g. valid poses of the egg such that it is inside an oven). These constraints can be readily integrated with existing constraints (e.g. avoiding collisions, respecting kinematics) within off-the-shelf TAMP systems. Thus, the overall system is able to generate solutions that not only respect constraints derived from the open world goal, but also are physically feasible on robot hardware.
@@ I Introduction
We propose OWL-TAMP (Open-World Language-based TAMP), an approach that integrates open world concepts via constraint generation into a TAMP system with traditional robotics operations and constraints. Our key contributions towards this framework are: (1) a method for generating constraints on action sequences to specify partial plans with language descriptions; (2) a method for generating constraints on continuous variables affected within the partial plan from (1); and (3) combining both (1) and (2) within a TAMP system. We test our framework on a variety of open world tasks in simulation and find that it is able to solve open world tasks with a higher success rate than several ablations, including a pure VLM or pure TAMP system respectively. We also demonstrate that OWL-TAMP enables a real-world robot to solve complex, long-horizon manipulation tasks specified through language directly from sensor input.
II Related Work
Task and Motion Planning (TAMP). A standard approach to solving long-horizon robot manipulation tasks via joint discrete-continuous reasoning is TAMP [11, 12]. Task planners can be made to solve open-world tasks expressed in natural language by translating the language into a symbolic goal expression conducive to planning [13]. However, this approach relies on the assumptions that: (1) the goal expression can be translated exactly into an expression over a set of predicates and (2) the translated expression can be achieved by planning with specific symbolic operators. Enabling TAMP systems to solve a wider range of tasks requires expanding their symbolic vocabulary either manually, or via learning [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By contrast, our approach leverages VLMs to directly write constraint functions that expand the set of tasks a TAMP system can solve without directly expanding its base symbolic vocabulary.
@@ II Related Work
LLMs and VLMs for Robotic Manipulation. Recent work has leveraged internet-scale foundation models to enable complex short-horizon visuomotor behaviors [26, 27, 4, 28, 29]. These approaches seek to solve relatively short-horizon tasks (such as pouring water into a cup or folding a t-shirt) whereas we are interested in longer horizon tasks composed of multiple atomic behaviors. Additionally, to the extent that these approaches do handle multi-step tasks (e.g. pick-and-place tasks), they often make assumptions or leverage heuristics specific to the particular class of tasks (e.g. that the task always consists of a series of picks followed by placements) [4], or require learning in simulation, and are thus not zero-shot [28, 29]. By contrast, our work proposes a clear contract for combining VLMs with generic domain-independent TAMP systems, with no requirement for learning in simulation, and minimal assumptions on the class of tasks that the system may be asked to perform.
LLMs and VLMs for Robot Task Planning. Another line of work has leveraged foundation models to solve long horizon planning problems. These works have focused on the problem of sequencing discrete skills in restricted settings where continuous parameters are either unnecessary, or can be easily determined by simple heuristics [5, 30, 31, 6, 32]. Some works have considered tasks where continuous parameters are important [32, 33, 34, 35, 36, 37, 38], but have relied on foundation models to either directly output such parameters, or to write code to do so. Thus, these approaches have generally only been able to handle problems in which the continuous parameters can be easily determined, most do not consider resolving complex constraints like reachability, kinematic feasibility or grasp stability. The few approaches that solve complex tasks that require precise discrete and continuous parameters [39, 40, 41, 42, 7] either have limited ability to handle cases where the system must consider alternative symbolic plans to achieve a goal (e.g. moving an object out of the way in order to place an object), or generally struggle with long-horizon tasks (where explicit search tends to perform better than LLMs [43, 44]). By contrast, our approach leverages off-the-shelf TAMP systems for planning and is thus able to solve longer horizon tasks with more complex constraints. Concurrent work [45] attempts to define a similar constraint-based contract between LLMs and TAMP systems for manipulation tasks, but focuses only on discrete constraints whereas our approach is able to leverage VLMs for both discrete and continuous constraint generation.
III Problem Setup
We adopt a model-based mixed discrete-continuous planning approach to controlling a robot to solve open-world tasks. We assume a planning model which contains commonplace manipulation primitives applicable across a very wide range of tasks and leverage foundation models to extend the core model to reason about novel, task-specific dynamics and constraints. Intuitively, the underlying planning model should capture generic dynamics and constraints (e.g. inverse kinematics and reachability, collision constraints) that apply across any task a robot might be faced with, while the foundation model should provide additional task-specific constraints (e.g. that an object must be placed in a pan for it to be ‘cooked’, that serving coffee in a mug requires that mug be upright) that serve to specialize the planning model to the given situation.
We model the system using a Planning Domain Definition Language (PDDL)-style [46] factored action language, which represents states and actions in terms of predicates. Our approach is not limited to this representational choice; it is compatible with multiple planning frameworks, e.g. PDDLStream [9] and SeSaME [12, 8, 16]. In PDDL, state variables are represented as literals, true or false evaluations of predicates for particular values of their parameters.
We use a single robot acting in a simplified manipulation domain as a pedagogical running example. Because robotics inherently involves continuous values, we consider not only discrete parameter types but also continuous ones, namely: obj - a discrete manipulable object $o$, conf - a continuous robot configuration $q\in\mathbf{R}^{d}$, traj - a continuous robot trajectory comprised of a sequence $n$ of configurations $\tau\in\mathbf{R}^{nd}$, grasp - a continuous object grasp pose $g\in\mathrm{SE}(3)$, and pose - a continuous object placement pose $p\in\mathrm{SE}(3)$.
The fluent predicates, i.e. predicates with truth values that can change over time, are: AtConf($q$: conf) - the robot is currently at configuration $q$, HandEmpty() - the robot’s hand is currently empty, AtPose($o$: obj, $p$: pose) - object $o$ is currently at placement pose $p$, and AtGrasp($o$: obj, $g$: grasp) - object $o$ is currently grasped with grasp pose $g$. From these predicates, we can describe states, which are represented by true literals. For example, the initial state in a domain with a single object apple might be: $s_{0}=[{\texttt{AtConf}}(q_{0}),{\texttt{HandEmpty}}(),{\texttt{AtPose}}({\texttt{apple}},p_{0}),...].$
Figure 1: Hybrid traditional and VLM constraint satisfaction. The constraint network for a candidate plan to pick and place the apple that achieves the goal “put the orange on the table where the apple initially is” in Figure LABEL:fig:system. Due to the Collision and VLMPose constraints, this constraint network isn’t satisfiable; however, a plan that first moves apple out of the way before placing apple is.
@@ III Problem Setup
Parameterized actions, which the robot can apply to affect a change in a state, are defined by a name, list of typed parameters, list of static literal constraints (con) that the parameters must satisfy, list of fluent literal preconditions (pre) that must hold before applying the action, and list of fluent literal effects (eff) that hold in the state after applying the action. The actions move and attach model the robot moving between two configurations and attaching an object to itself, for example, by grasping it.
move$(q_{1}:\text{conf},\;q_{2}:\text{conf},\;\tau:\text{traj})$
con: [Motion$(q_{1},\tau,q_{2})$]
pre: [AtConf$(q_{1})$]
eff: [AtConf$(q_{2})$, $\neg$¬AtConf$(q_{1})$]
attach$(o:\text{obj},\;p:\text{pose},\;g:\text{grasp},\;q:\text{conf})$
con: [Kin$(q,o,g,p)$]
pre: [AtPose$(o,p)$, HandEmpty()()( ), AtConf$(q)$]
eff: [AtGrasp$(o,g)$, $\neg$¬AtPose$(o,p)$, $\neg$¬HandEmpty()()( )]
Ground action instances of these parameterized actions must satisfy the following static predicates: Motion($q_{1}$: conf, $\tau$: traj, $q_{2}$: conf) - $\tau$ is a valid trajectory that connects configurations $q_{1}$ and $q_{2}$, and Kin($q$: conf, $o$: obj, $g$: grasp, $p$: placement) - configuration $q$ satisfies a kinematics constraint with placement pose $p$ when object $o$ is grasped with grasp pose $g$.
III-A Open World Predicates and Actions
So far, we’ve described a small and finite set of traditional TAMP predicates and actions. Intuitively, these correspond to generic dynamics and constraints that a robot encounters due to its embodiment in the physical world. However, we are interested in modeling and planning with open-world concepts that are environment or task specific. To do so, we propose parameterizing select predicates and actions with an additional type, a description $d$. Descriptions modify the semantics of predicates and actions to respect an open-world natural-language instruction. Descriptions help specialize the overly general robot interactions (e.g. moving without collision, grasping stably) in the traditional planning model to achieve novel outcomes. Overall, this strategy can be seen as bootstrapping an unbounded set of predicates and actions from a finite set by leveraging language itself as a parameter.
Consider the VLMPose($d$: description, $o$: obj, $p$: pose) constraint, which is true if object $o$ at placement $p$ satisfies description $d$. Some example descriptions $d$ are: “orange at the center of the table”, “orange at the apple’s initial location”, and “orange as far way from the robot as possible”. Using this constraint, we can formulate a detach action, which involves the robot releasing object $o$ according to the description $d$. This can correspond to placing the object on a surface, stacking the object on another object, dropping the object in a bin, inserting the object into an outlet, etc.
detach$(d:\text{description},\;o:\text{obj},\;g:\text{grasp},\;p:\text{pose},\;q:\text{conf})$
con: [Kin$(q,o,g,p)$, VLMPose$(d,o,p)$]
pre: [AtPose$(o,p)$, HandEmpty()()( ), AtConf$(q)$,
eff: [AtGrasp$(o,g)$, $\neg$¬AtPose$(o,p)$, $\neg$¬HandEmpty()()( )]
We can define additional parameterized actions that model different interaction types, such as an action that moves a cup through waypoints to fill it up or pour out of it  [^1] We use demonstrate this in the ‘Mug3’ and ‘SoupPour’ tasks in Section V..
A key advantage of our approach is that we plan with both traditional robot constraints as well as task-specific open-world constraints. Consider the problem in Figure LABEL:fig:system, where the the goal is to ‘put the orange on the table where the apple initially is”. Figure 1 (left) displays the simplified constraint network [11], a bipartite graph from free action parameters (in bold) to the action constraints they are involved in (conf), induced by a plan that directly picks and places the apple:
$\pi$ $=[...,{\texttt{attach}}({\texttt{apple}},p_{0}^{A},\bm{g},\bm{q_{1}}),...,{\texttt{detach}}($
[…]
This constraint network is unsatisfiable because the VLMPose constraint restricts the set of placements that satisfy the task and the Collision constraint prevents unsafe placements. But through the use of TAMP, our approach can backtrack over candidate plans that first move the apple to eventually find a satisfiable constraint network and ultimately a solution.
IV TAMP with Open World Concepts
We address TAMP problems $\langle s_{0},{\cal A},g\rangle$ described by an initial state $s_{0}$, set of parameterized actions ${\cal A}$, and goal $g$. Unlike traditional TAMP problems, the goal $g$ is not a logical formula over literals but rather is a goal description provided in English text. Thus, solving such problems requires translating $g$ into some form that can be used within TAMP.
One approach to this translation would be to directly prompt a VLM to output some logical formula over literals (which we will denote as $G$) from the goal description $g$ [13]. Given this, one could simply call an off-the-shelf TAMP system to achieve $G$. While this approach is straightforward, and powerful, it is limited in the kinds of tasks it is able to express in at least two ways: (1) it can only define a goal state to achieve and cannot specify intermediate behaviors or states that need to occur before the goal, and (2) it can only express goals in terms of predicates that are already built into the TAMP system.
[…]
We seek to address these limitations in the expressivity of direct translation by instead translating $g$ into more flexible discrete and continuous constraints (as depicted in Figure LABEL:fig:system). Specifically, we first prompt a VLM to supply a set of discrete constraints over open world action orderings, and then induce continuous constraints in the form of code for particular predicates (such as VLMPose) that appear in the effects or constraints of action definitions used as part of our first stage. We then incorporate these constraints into a TAMP system such that it only yields plans that satisfy these constraints. Intuitively, these constraints will be task specific and enable the system to achieve tasks it otherwise could not. Conversely, through using a TAMP system, OWL-TAMP inherits theoretical guarantees with respect to the non-VLM constraints such as plan soundness, which is critical for safety, and probabilistic completeness [47]. In the cooking task mentioned above, generating a discrete constraint that any valid plan should execute a detach(strawberry, pan) action before a detach(strawberry, bowl) action would be sufficient to enable the TAMP system to solve the task. Similarly, in the fruit sorting task, all that is required is a continuous constraint on the outcome of every detach(fruit) for a TAMP system to accomplish the underlying goal.
In what follows, we first describe our procedure for discrete constraint generation before describing how we generate continuous constraints given initial discrete constraints.
IV-A Generating discrete planning constraints with a VLM
Given a goal description $g$, we first prompt a VLM to generate a partial plan that serves as a discrete constraint on the space of TAMP solutions. To enable this, we associate a natural language description of each available action with that particular action. Although we could directly prompt a VLM for relevant actions and goals, without a list of candidates, the VLM is likely to be syntactically and semantically inaccurate. Instead, we propose to first ground the set of reachable actions $A$ and literals $L$ available to the TAMP system before prompting the VLM to return values in these sets. We use relaxed planning [48, 49] from the initial state $s_{0}$ to simultaneously ground [50, 51, 52] and explore the sets of reachable actions $A$ and literals $L$. When instantiating continuous parameters, we use placeholder values, such as optimistic values [53, 47, 9, 54], to ensure a finite set of actions are instantiated. Similarly, we use placeholders for description parameters.
Algorithm 1 VLM Task Reasoning
1:procedure vlm-task-reasoning($s_{0},{\cal A},g$)
2:    $A\leftarrow\textsc{ground-actions}(s_{0},{\cal A})$
[…]
4:    $[a_{1},...,a_{n},l_{1},..,l_{m}]\leftarrow\textsc{query-vlm}(\text{``What partial plan}\newline \text{using actions }\{A\}\text{ for goal literals }\{L\}\text{ achieves goal }\{g\}\text{?''})$” )
5:    for $i\in[1,n-1]$ do
[…]
9:    $G\leftarrow\{l_{1},..,l_{m}\}$
[…]
Algorithm 1 presents the VLM partial plan generation pseudocode. It takes in a TAMP problem $\langle s_{0},{\cal A},g\rangle$, where $g$ is a text goal description. It first grounds the set of actions $A$ reachable from $s_{0}$ using ground-actions. Then, it accumulates the set of reachable literals $L$ by taking the effects of all actions $A$. These sets can be filtered by action or predicate type if it is desired to focus VLM assistance on specific aspects of the planning problem. Then, it prompts query-vlm for a partial plan $[a_{1},...,a_{n},l_{1},...,l_{k}]$ using actions $a_{i}\in A$ and goal literals $l_{j}\in L_{m}$ that achieve the goal description $g$. Importantly, we have the VLM fill in the description parameter $d$ for each of these actions. We then transform the original TAMP problem to force solutions to admit the partial plan as a subsequence. Specifically, we create a predicate Executed that models whether the $i$th action in the plan was executed and add Executed to the effects of action $a_{i}$ and the preconditions of action $a_{i+1}$. Finally, we make the planning goal be $G=\{l_{i},...,l_{m}\}\subseteq L$ and $\textsc{Executed}(n)$, which indicates that all actions have been executed and solve the transformed TAMP problem with a generic TAMP algorithm.
Consider the cooking problem mentioned earlier in the beginning of Section IV $g=$“Cook the strawberry by putting it in the pan, then finally serve it in the bowl”. Suppose the VLM returns no goal literals, but just the partial plan:
@@ IV-A Generating discrete planning constraints with a VLM
Although the VLM plan $\vec{\pi}$ does capture the intent of the task (i.e., to place the apple in the pan before serving it), this plan is not legal because objects must be picked with the attach action before they can be detached. Fortunately, the underlying TAMP system models this, and thus providing this partial plan, along with the generated Executed predicates, to the TAMP system will result in the TAMP system generating legal plans that are at least 4 actions long.
Figure 2: Simulated RAVENS-YCB Manipulation Tasks. Intermediate states during execution of our approach on 3 simulated tasks.
Figure 3: Real-World Tasks. Two tasks from our real-world system deployment (Section V) and their corresponding goal descriptions.
IV-B Grounding Continuous Constraints with a VLM
The procedure in Section IV-A enables us to generate actions with language parameters (Section III-A) fully specified. However, in order to correctly apply these actions, we need to be able to interpret how the language parameter should affect legal action parameter values. More specifically, we need to provide an implementation for any constraint fluents (such as the VLMPose$(d,o,p)$ fluent introduced in Section III-A) that use the language description $d$.
More specifically, consider the coffee task (i.e. where $g=$ “Can you setup the cup on the table so I can properly pour coffee into it?”), and suppose the discrete generation procedure has produced a plan that contains the following action:
@@ IV-B Grounding Continuous Constraints with a VLM
ensuring it is upright and positioned

### v4

# Open-World Task and Motion Planning via Vision-Language Model Generated Constraints

@@ Abstract
Foundation models like Vision-Language Models (VLMs) excel at common sense vision and language tasks such as visual question answering. However, they cannot yet directly solve complex, long-horizon robot manipulation problems requiring precise continuous reasoning. Task and Motion Planning (TAMP) systems can handle long-horizon reasoning through discrete-continuous hybrid search over parameterized skills, but rely on detailed environment models and cannot interpret novel human objectives, such as arbitrary natural language goals. We propose integrating VLMs into TAMP systems by having them generate discrete and continuous language-parameterized constraints that enable open-world reasoning. Specifically, we use VLMs to generate discrete action ordering constraints that constrain TAMP’s search over action sequences, and continuous constraints in the form of code that augments traditional TAMP manipulation constraints. Experiments show that our approach, OWL-TAMP, outperforms baselines relying solely on TAMP or VLMs across several long-horizon manipulation tasks specified directly in natural language. We additionally demonstrate that OWL-TAMP can be deployed with an off-the-shelf TAMP system to solve challenging manipulation tasks on real-world hardware.
1 Introduction
Solving complex, long-horizon manipulation tasks is a fundamental challenge in robotics. Recently, there has been significant interest in solving such tasks directly from language and image input [27, 10, 36]. In this setting, a user provides a command like “put away all the cutlery in the utensil holder” and the robot must execute a sequence of actions to achieve the intended effect. Even with pre-engineered manipulation controller primitives, such as those for moving, picking objects, and placing objects, these problems remain difficult: the robot must correctly interpret the command and sequence actions with the correct discrete arguments, such as which objects to grasp, and continuous arguments, such as grasps, stable placement poses and collision-free motions.
Consider the task in Figure 1, where the robot must “put the banana near where the other fruits are initially.” One approach is to prompt a Vision-Language Model (VLM) [1] with the initial image, task, and available controllers [25]. VLMs excel at interpreting commands and leveraging common sense, and would likely suggest picking the banana before placing it. However, they struggle to reason about continuous parameter values and geometry (e.g., synthesizing collision-free grasps and placements) without task-specific fine-tuning [5, 10]. Here, the robot’s gripper cannot safely grasp the banana without first moving obstructing objects. By contrast, classical Task and Motion Planning (TAMP) systems explicitly reason over discrete and continuous values using symbolic predicates, operators, and geometric models [46, 14]. A TAMP system would detect the blocked grasp and plan to move the milk carton before picking the banana. However, TAMP systems require extensive manual modeling and cannot handle novel concepts without additional engineering. In this case, the TAMP system lacks the concept for “near”, and thus cannot place the banana correctly “near” the other fruits.
In this work, we combine the strengths of foundation models and TAMP to solve manipulation tasks specified in natural language. Our key insight is that we can leverage the commonsense and language-grounding capabilities of VLMs to guide TAMP systems through dynamically generated constraints. Specifically, VLMs map language into both code expressing continuous action constraints (e.g., valid poses “near” other fruits) and discrete constraints plan sketches (e.g., pick the banana, then place it “near” the other fruits). These constraints are combined with common, pre-existing TAMP constraints like collision avoidance and kinematics, and are used by a TAMP system to solve for a plan that jointly satisfies all constraints. In our example, the TAMP system produces a correct plan that moves the milk carton to enable grasping the banana, then places it near the other fruit as dictated by the VLM-generated continuous pose constraint. By integrating such constraints, TAMP systems become “open-world” – capable of solving tasks beyond their pre-defined symbolic vocabulary (e.g., reasoning about “near” or “oriented straight” without explicit pre-defined predicates). As a result, our approach is able to solve tasks beyond the closed-world scope of prior work that just use VLMs to accelerate discrete planning [51].
@@ 1 Introduction
We propose OWL-TAMP (Open-World Language-based TAMP), our approach that integrates open-world concepts into TAMP via constraint generation over actions and continuous variables. Our main contribution is to propose a clear contract — in the form of constraints — for combining VLMs with generic domain-independent TAMP systems to enable zero-shot generalization to a variety of tasks. More specifically, OWL-TAMP: (1) generates constraints on action sequences from language to specify plan sketches; (2) generates constraints on continuous variables within these plans via code; and (3) integrates both into standard TAMP pipelines. In simulation, OWL-TAMP achieves higher success rates on open-world tasks than several baselines, including pure VLM and pure TAMP approaches. We also demonstrate that it enables a real robot to solve complex, natural-language based long-horizon tasks with initial states given as raw camera images.
Figure 1: OWL-TAMP Overview. OWL-TAMP uses VLMs to generate task-specific open-world constraints that augment traditional robot constraints like kinematics and grasp stability. The VLM first generates (a) plan sketches, open-world actions with natural language descriptions, and uses these to generate (b) continuous constraints in the form of code that check whether actions achieve their intended open-world predicate effects. (c) A TAMP system consumes these constraints to solve for a plan that satisfies both traditional and VLM-generated constraints. Here, OWL-TAMP moves the milk carton, enabling the banana to be grasped successfully.
2 Related Work
Task and Motion Planning (TAMP). A standard approach for solving long-horizon robot tasks via joint discrete-continuous reasoning is TAMP [16, 35]. TAMP can solve tasks expressed in natural language by translating language into symbolic goals [50], but assumes (1) exact translation into a pre-defined predicate set and (2) that these goals are achievable with existing symbolic operators. For instance, solving tasks like Figure 1 would require a “near” predicate; if only an “on top” predicate exists, translation would fail. Expanding TAMP’s capabilities thus requires manual or learned predicate invention [33, 43, 37, 3]. In contrast, our approach uses pretrained VLMs without finetuning to generate constraints that expand TAMP’s capabilities without task-specific demonstration or exploration data (Section 6).
@@ 2 Related Work
LLMs, VLMs and VLAs for Robotic Manipulation. Recent work has leveraged VLMs [28, 26, 27, 40] to enable complex short-horizon visuomotor behaviors, such as pouring water into a cup. In contrast, we focus on longer horizon tasks composed of multiple atomic behaviors. Additionally, to the extent these approaches do handle multi-step tasks (e.g., pick-and-place tasks), they often make assumptions or leverage heuristics specific to the particular class of tasks — such as hardcoding the action ordering [27] or learning in simulation — and are thus not zero-shot [40]. More recent works attempt to solve long-horizon tasks by training Vision Language Action models (VLAs) on large amounts of robot demonstration data [5, 42]. However, these models require finetuning on some task and environment-specific data when adapting to novel setups and embodiments. By contrast, our approach combines VLMs with generic, domain-independent TAMP systems to solve tasks with no requirement for additional demonstrations or learning in simulation.
LLMs and VLMs for Robot Task Planning. Another line of work uses foundation models for planning by sequencing discrete skills in settings where continuous parameters are either unnecessary or easily determined by heuristics [31, 29, 44, 30]. Some works address continuous parameters [30, 22, 2, 36, 48, 52, 12], but rely on foundation models to predict them directly or generate code, typically succeeding only when parameters are easily inferred from images or text (e.g., top-down grasping). A few approaches tackle complex tasks requiring precise discrete and continuous parameters [45, 11, 8, 10], but struggle with discrete-level backtracking over action sequences or long-horizon tasks where explicit search is preferable [32]. Our method instead combines VLMs with off-the-shelf TAMP systems to handle long-horizons and more complex constraints. Recent work [51] leverages VLMs to generate subgoals for a TAMP system to solve sequentially, but does not generate any continuous constraints. Concurrent work [21] also defines a constraint-based contract between LLMs and TAMP, but only handles discrete constraints and requires a custom TAMP solver.
Overall, while there are prior works that explore the general idea of combining foundation models with search-based planning, our approach is the first to directly integrate with full off-the-shelf TAMP systems. In OWL-TAMP, the foundation model is not simply accelerating search but rather writing open-world continuous and discrete constraints to augment the model that search is being performed on.
3 Problem Formulation
We aim to solve tasks specified in natural language, given an initial scene observation, a set of objects, a library of manipulation controllers, and a pre-existing Task and Motion Planning (TAMP) system with symbolic components. We define a task as a tuple $\langle s_{0},\mathcal{A},\mathcal{I},L\rangle$, where $s_{0}$ is the initial state of the robot and a set of named objects (identified via object detection and segmentation) in the environment (e.g., joint positions, object poses obtained from vision-based pose estimation), $\mathcal{A}$ is a set of parameterized actions, $\mathcal{I}$ is a set of initial image observations of the scene, and $L$ is the natural language instruction (e.g., “put the banana near the other fruit” in Figure 1).
Each parameterized action in $\mathcal{A}$ consists of a list of typed parameters, which may be discrete (e.g., object identifiers) or continuous (e.g., object poses or robot configurations), and a low-level controller (e.g., a joint-trajectory controller) that executes the action according to the parameter values. For example, consider the typed parameters of an attach action: attach($o{:}\;{\texttt{obj}},p{:}\;{\texttt{pose}},g{:}\;{\texttt{grasp}},q{:}\;{\texttt{conf}}$). Here, $o$ is a discrete object identifier, $p\in\text{SE}(3)$ is an approach pose near object $o$, $g\in\text{SE}(3)$ is a grasp pose where the robot closes its fingers, and $q\in\mathbb{R}^{d}$ is the corresponding robot joint configuration used to execute grasp $g$ on $o$. Specifying concrete values for all parameters grounds the action, yielding a fully executable manipulation controller.
The robot’s objective is to generate and execute a plan: a sequence of grounded actions which, when applied from the initial state $s_{0}$, results in a final state $s_{*}$ that satisfies the goal described by $L$.
3.1 Task and Motion Planning
Solving the tasks described above is challenging because the robot must not only select a valid sequence of actions, but also ground each action with precise parameters that achieve the goal described by $L$. We are interested in leveraging a TAMP system to help address these challenges; we thus assume access to a model-based TAMP system that is fixed and shared across all tasks.
Figure 2: Real-World Tasks. OWL-TAMP (a) reasons to move the milk obstructing the banana, then places the banana near the other fruit (Figure 1). (b) plans placements satisfying both the language goal and collision constraints. (c) identifies the block as shortest, and weighs it on the scale before disposing.
Within TAMP, we model states and actions using a Planning Domain Definition Language (PDDL)-style [41] factored action language, where states are sets of true-valued literals, i.e., predicates applied to specific parameters. For example, the initial state $s_{0}$ in Figure 1 includes [AtConf($q_{0}$), HandEmpty(), AtPose(banana, $p_{0}$), On(banana, table), …], indicating that the robot is at configuration $q_{0}$, its gripper is empty, the banana is at placement pose $p_{0}$ and on the table.
@@ 3.1 Task and Motion Planning
Each parameterized action $a$ in $\mathcal{A}$, defined as part of the task, is extended in the TAMP system with a set of constraints (con) that the parameters must satisfy, preconditions (pre) that must hold before applying the action, and effects (eff) that describe how the state changes after executing action $a$. The actions move and attach model the robot moving between two configurations and attaching an object to itself via grasping, respectively.
@@
con: [Kin(q,o,g,p)(q,o,g,p)]
pre: [AtPose(o,p)(o,p), HandEmpty()(), AtConf(q)(q)]